A spinning wheel has a moment of inertia of 2.50kg.m^2 about its axis of rotation starting from rest,how much constant torque is required to bring it to an regular speed of 300 rpm in 10.0s?.

1
Expert's answer
2022-01-26T17:45:44-0500

1) Find the angular acceleration:


"\alpha=\omega/t."


2) Find the torque:


"\tau=I\alpha=I\omega/t=\pi In/(30t)=7.85\text{ N}·\text m."